Seismic activity over the last 24 hours has been similar to that observed over the previous few days. Sixteen long-period earthquakes were recorded by most of the stations in the network, along with several signals generated by rockfalls on the dome. The moderate-amplitude broadband tremor recorded only at the Gages seismic station has continued. Two small volcano-tectonic earthquakes were observed, with locations at 4 km beneath Long Ground and 2 km beneath Harris's.
No EDM measurements were possible today because of the persistent low cloud over the volcano. When the cloud lifted slightly in the late afternoon, an observation flight was made in the helicopter. Most of the crater area was obscured by cloud, but the scientists noted several new features in the eastern side of the dome. No rockfalls were observed, but the dome continues to steam vigorously.